OpenStudy (anonymous):

limit as x goes to 0: (1-sec^2(2x))/((6x)^2)

OpenStudy (anonymous):

0/0 L'Hopital

OpenStudy (anonymous):

does that mean DNE? how come?

OpenStudy (anonymous):

Im going to teach myself l'hopital tonight

OpenStudy (anonymous):

it isn't DNE though - was wrong answer?

OpenStudy (anonymous):

If you are very early in calculus, you can get away with saying dne, but soon you will be taught how to manipulate it and get good answer.

OpenStudy (anonymous):

DNE was the marked incorrect

OpenStudy (anonymous):

I didn't say it was 'dne'

OpenStudy (anonymous):

haha: :-)

OpenStudy (anonymous):

im trying to calc it without using l'hopital

OpenStudy (anonymous):

You can try and manipulate in a way that the limit does not go to 0/0

OpenStudy (anonymous):

was going to change it to sin^2(2x)

OpenStudy (anonymous):

It still gives you 0/0. Take derivative of top, and take derivative of bottom

myininaya (myininaya):

\[\frac{\tan^2(2x)}{(6x)^2}=\frac{\frac{\sin^2(2x)}{\cos^2(2x)}}{(6x)^2}=\frac{\sin^2(2x)}{(6x)^2\cos^2(2x)}\] \[=\frac{\sin^2(2x)}{36x^2\cos^2(2x)}=\frac{1}{36}*4*\frac{\sin(2x)\sin(2x)}{(2x)(2x)(\cos^2(2x))}\]

myininaya (myininaya):

1/36*4*1*1*1=4/36=1/9

OpenStudy (anonymous):

why do the sin and cos cross out?

myininaya (myininaya):

sin2x/2x->1 as x->0 sin2x/2x->1 as x->0 1/cos^(2x)->1/1=1 as x->0

OpenStudy (anonymous):

oh, the rules!

OpenStudy (anonymous):

thank you! Myininaya!

myininaya (myininaya):

i answered that previous one with l'hospital if you want to look at it it is an attachment

myininaya (myininaya):

without*

OpenStudy (anonymous):

oh sweet! thanks!

myininaya (myininaya):

lol i keep saying with i mean without lol

myininaya (myininaya):

thanks chag and alkileez for medals lol

OpenStudy (anonymous):

Good work. If you haven't got to L'hopital yet, you have to get creative like this.